OpenStudy (anonymous):

Polynomials How many Polynomials (with real coefficients) are there of degree 10 such that \[P(x^2)=P(x) .P(x-1)\]

OpenStudy (anonymous):

lets think together... i dont have any idea

OpenStudy (anonymous):

i think.....in p(x^2) the exponents of x are even.... but but in p(x)P(x-1) the exponents of x may be even or odd unless a specific condition is present.... and i have no clue about this condition!!!

OpenStudy (anonymous):

i found this hint Note first that p(1) = p(1) p(0). So p(1) = 0 or p(0) = 1. Suppose that p(1) = 0. Then p(2^2) = p(2) p(1) = 0. Continuing in this way, we deduce that p(n) = 0 for infinitely many positive integers n. As a nonconstant polynomial, p(x) can have only finitely many roots.Thus we have a contradiction, and can conclude that p(0) = 1. Suppose p(a) = 0 for some complex number a. Then p(a^2) = p(a^4) =· · · = 0. However, a polynomial can only have a finite number of roots. So a must be a root of unity. Therefore, |a| = 1. Similarly, we can prove that |a + 1| = 1.

OpenStudy (anonymous):

now if we let a=x+iy with some simple calculations we get \[a=-\frac{1}{2}\pm i \frac{\sqrt{3}}{2}\] since |a| = 1 and |a + 1| = 1

OpenStudy (anonymous):

then they are the only roots of a such polynomial now i think we can say \[P(x)=c((x+\frac{1}{2}-i \frac{\sqrt{3}}{2})(x+\frac{1}{2}+i \frac{\sqrt{3}}{2}))^n=c(x^2+x+1)^n\]

OpenStudy (anonymous):

P(0)=1 then c=1 and 2n=10 --> n=5 so i think the only answer is \[P(x)=(x^2+x+1)^5\]

OpenStudy (anonymous):

@Zarkon @experimentX

OpenStudy (anonymous):

@satellite73 plz check the solution if u can

OpenStudy (anonymous):

That is what I got \[ \left(x^2+x+1\right)^5 \]

OpenStudy (anonymous):

I solve the following equations involving the coefficient using Mathematica \[ a(0) (a(0)-a(1)+a(2)-a(3)+a(4)-a(5)+a(6)-a(7)+ a(8)-a(9)+1)-a(0)=0\\ 2 a(9)-10=0\\ 2 a(7)-8 a(8)+a(8) (a(9)-10)+(a(8)-9 a(9)+45) a(9)+36 a(9)-120=0\\ 2 a(8)+(a(9)-10) a(9)-10 a(9)+45=0\\ a(1) (a(0)-a(1)+a(2)-a(3)+a(4)-a(5)+a(6)-a(7)+ a(8)-a(9)+1)+a(0) (a(1)-2 a(2)+3 a(3)-4 a(4)+5 a(5)-6 a(6)+7 a(7)-8 a(8)+9 a(9)-10)=0\\ (a(1)-2 a(2)+3 a(3)-4 a(4)+5 a(5)-6 a(6)+7 a(7)-8 a(8)+9 a(9)-10) a(1)-a(1)+a(0) (a(2)-3 a(3)+6 a(4)-10 a(5)+15 a(6)-21 a(7)+28 a(8)-36 a(9)+45)+a(2) (a(0)-a(1)+a(2)-a(3)+a(4)-a(5)+a(6)-a(7)+ a(8)-a(9)+1)=0\\ 2 a(5)-6 a(6)+21 a(7)-56 a(8)+a(7) (a(8)-9 a(9)+45)+a(6) (a(9)-10)+(a(6)-7 a(7)+28 a(8)-84 a(9)+210) a(9)+126 a(9)+a(8) (a(7)-8 a(8)+36 a(9)-120)-252=0\\ 2 a(6)-7 a(7)+27 a(8)+a(8) (a(8)-9 a(9)+45)+a(7) (a(9)-10)-84 a(9)+a(9) (a(7)-8 a(8)+36 a(9)-120)+210=0\\ a(1) (a(2)-3 a(3)+6 a(4)-10 a(5)+15 a(6)-21 a(7)+28 a(8)-36 a(9)+45)+a(3) (a(0)-a(1)+a(2)-a(3)+a(4)-a(5)+a(6)-a(7)+ a(8)-a(9)+1)+a(2) (a(1)-2 a(2)+3 a(3)-4 a(4)+5 a(5)-6 a(6)+7 a(7)-8 a(8)+9 a(9)-10)+a(0) (a(3)-4 a(4)+10 a(5)-20 a(6)+35 a(7)-56 a(8)+84 a(9)-120)=0\\ (a(2)-3 a(3)+6 a(4)-10 a(5)+15 a(6)-21 a(7)+28 a(8)-36 a(9)+45) a(2)-a(2)+a(0) (a(4)-5 a(5)+15 a(6)-35 a(7)+70 a(8)-126 a(9)+210)+a(4) (a(0)-a(1)+a(2)-a(3)+a(4)-a(5)+a(6)-a(7)+ a(8)-a(9)+1)+a(3) (a(1)-2 a(2)+3 a(3)-4 a(4)+5 a(5)-6 a(6)+7 a(7)-8 a(8)+9 a(9)-10)+a(1) (a(3)-4 a(4)+10 a(5)-20 a(6)+35 a(7)-56 a(8)+84 a(9)-120)=0\\ a(1) (a(4)-5 a(5)+15 a(6)-35 a(7)+70 a(8)-126 a(9)+210)+a(3) (a(2)-3 a(3)+6 a(4)-10 a(5)+15 a(6)-21 a(7)+28 a(8)-36 a(9)+45)+a(5) (a(0)-a(1)+a(2)-a(3)+a(4)-a(5)+a(6)-a(7)+ a(8)-a(9)+1)+a(4) (a(1)-2 a(2)+3 a(3)-4 a(4)+5 a(5)-6 a(6)+7 a(7)-8 a(8)+9 a(9)-10)+a(2) (a(3)-4 a(4)+10 a(5)-20 a(6)+35 a(7)-56 a(8)+84 a(9)-120)+a(0) (a(5)-6 a(6)+21 a(7)-56 a(8)+126 a(9)-252)=0\\ (a(3)-4 a(4)+10 a(5)-20 a(6)+35 a(7)-56 a(8)+84 a(9)-120) a(3)-a(3)+a(2) (a(4)-5 a(5)+15 a(6)-35 a(7)+70 a(8)-126 a(9)+210)+a(0) (a(6)-7 a(7)+28 a(8)-84 a(9)+210)+a(4) (a(2)-3 a(3)+6 a(4)-10 a(5)+15 a(6)-21 a(7)+28 a(8)-36 a(9)+45)+a(6) (a(0)-a(1)+a(2)-a(3)+a(4)-a(5)+a(6)-a(7)+ a(8)-a(9)+1)+a(5) (a(1)-2 a(2)+3 a(3)-4 a(4)+5 a(5)-6 a(6)+7 a(7)-8 a(8)+9 a(9)-10)+a(1) (a(5)-6 a(6)+21 a(7)-56 a(8)+126 a(9)-252)=0\\ a(3) (a(4)-5 a(5)+15 a(6)-35 a(7)+70 a(8)-126 a(9)+210)+a(1) (a(6)-7 a(7)+28 a(8)-84 a(9)+210)+a(5) (a(2)-3 a(3)+6 a(4)-10 a(5)+15 a(6)-21 a(7)+28 a(8)-36 a(9)+45)+a(7) (a(0)-a(1)+a(2)-a(3)+a(4)-a(5)+a(6)-a(7)+ a(8)-a(9)+1)+a(6) (a(1)-2 a(2)+3 a(3)-4 a(4)+5 a(5)-6 a(6)+7 a(7)-8 a(8)+9 a(9)-10)+a(0) (a(7)-8 a(8)+36 a(9)-120)+a(4) (a(3)-4 a(4)+10 a(5)-20 a(6)+35 a(7)-56 a(8)+84 a(9)-120)+a(2) (a(5)-6 a(6)+21 a(7)-56 a(8)+126 a(9)-252)=0\\ (a(4)-5 a(5)+15 a(6)-35 a(7)+70 a(8)-126 a(9)+210) a(4)-a(4)+a(2) (a(6)-7 a(7)+28 a(8)-84 a(9)+210)+a(6) (a(2)-3 a(3)+6 a(4)-10 a(5)+15 a(6)-21 a(7)+28 a(8)-36 a(9)+45)+a(0) (a(8)-9 a(9)+45)+a(8) (a(0)-a(1)+a(2)-a(3)+a(4)-a(5)+a(6)-a(7)+ a(8)-a(9)+1)+a(7) (a(1)-2 a(2)+3 a(3)-4 a(4)+5 a(5)-6 a(6)+7 a(7)-8 a(8)+9 a(9)-10)+a(1) (a(7)-8 a(8)+36 a(9)-120)+a(5) (a(3)-4 a(4)+10 a(5)-20 a(6)+35 a(7)-56 a(8)+84 a(9)-120)+a(3) (a(5)-6 a(6)+21 a(7)-56 a(8)+126 a(9)-252)=0\\ a(5) (a(4)-5 a(5)+15 a(6)-35 a(7)+70 a(8)-126 a(9)+210)+a(3) (a(6)-7 a(7)+28 a(8)-84 a(9)+210)+a(7) (a(2)-3 a(3)+6 a(4)-10 a(5)+15 a(6)-21 a(7)+28 a(8)-36 a(9)+45)+a(1) (a(8)-9 a(9)+45)+a(0) (a(9)-10)+(a(0)-a(1)+a(2)-a(3)+a(4)-a(5)+ a(6)-a(7)+a(8)-a(9)+1) a(9)+a(8) (a(1)-2 a(2)+3 a(3)-4 a(4)+5 a(5)-6 a(6)+7 a(7)-8 a(8)+9 a(9)-10)+a(2) (a(7)-8 a(8)+36 a(9)-120)+a(6) (a(3)-4 a(4)+10 a(5)-20 a(6)+35 a(7)-56 a(8)+84 a(9)-120)+a(4) (a(5)-6 a(6)+21 a(7)-56 a(8)+126 a(9)-252)=0\\ 2 a(0)-a(1)+a(2)-a(3)+a(4)-2 a(5)+a(6)-a(7)+a(8)+a(6) (a(4)-5 a(5)+15 a(6)-35 a(7)+70 a(8)-126 a(9)+210)+a(4) (a(6)-7 a(7)+28 a(8)-84 a(9)+210)+a(8) (a(2)-3 a(3)+6 a(4)-10 a(5)+15 a(6)-21 a(7)+28 a(8)-36 a(9)+45)+a(2) (a(8)-9 a(9)+45)+a(1) (a(9)-10)-a(9)+a(9) (a(1)-2 a(2)+3 a(3)-4 a(4)+5 a(5)-6 a(6)+7 a(7)-8 a(8)+9 a(9)-10)+a(3) (a(7)-8 a(8)+36 a(9)-120)+a(7) (a(3)-4 a(4)+10 a(5)-20 a(6)+35 a(7)-56 a(8)+84 a(9)-120)+a(5) (a(5)-6 a(6)+21 a(7)-56 a(8)+126 a(9)-252)+1=0\\ 2 a(1)-2 a(2)+3 a(3)-4 a(4)+5 a(5)-6 a(6)+7 a(7)-8 a(8)+a(7) (a(4)-5 a(5)+15 a(6)-35 a(7)+70 a(8)-126 a(9)+210)+a(5) (a(6)-7 a(7)+28 a(8)-84 a(9)+210)+a(3) (a(8)-9 a(9)+45)+a(2) (a(9)-10)+(a(2)-3 a(3)+6 a(4)-10 a(5)+15 a(6)-21 a(7)+28 a(8)-36 a(9)+45) a(9)+9 a(9)+a(4) (a(7)-8 a(8)+36 a(9)-120)+a(8) (a(3)-4 a(4)+10 a(5)-20 a(6)+35 a(7)-56 a(8)+84 a(9)-120)+a(6) (a(5)-6 a(6)+21 a(7)-56 a(8)+126 a(9)-252)-10=0\\ 2 a(2)-3 a(3)+6 a(4)-10 a(5)+14 a(6)-21 a(7)+28 a(8)+a(8) (a(4)-5 a(5)+15 a(6)-35 a(7)+70 a(8)-126 a(9)+210)+a(6) (a(6)-7 a(7)+28 a(8)-84 a(9)+210)+a(4) (a(8)-9 a(9)+45)+a(3) (a(9)-10)-36 a(9)+a(5) (a(7)-8 a(8)+36 a(9)-120)+a(9) (a(3)-4 a(4)+10 a(5)-20 a(6)+35 a(7)-56 a(8)+84 a(9)-120)+a(7) (a(5)-6 a(6)+21 a(7)-56 a(8)+126 a(9)-252)+45=0\\ 2 a(3)-4 a(4)+10 a(5)-20 a(6)+35 a(7)-56 a(8)+a(7) (a(6)-7 a(7)+28 a(8)-84 a(9)+210)+a(5) (a(8)-9 a(9)+45)+a(4) (a(9)-10)+(a(4)-5 a(5)+15 a(6)-35 a(7)+70 a(8)-126 a(9)+210) a(9)+84 a(9)+a(6) (a(7)-8 a(8)+36 a(9)-120)+a(8) (a(5)-6 a(6)+21 a(7)-56 a(8)+126 a(9)-252)-120=0\\ 2 a(4)-5 a(5)+15 a(6)-36 a(7)+70 a(8)+a(8) (a(6)-7 a(7)+28 a(8)-84 a(9)+210)+a(6) (a(8)-9 a(9)+45)+a(5) (a(9)-10)-126 a(9)+a(7) (a(7)-8 a(8)+36 a(9)-120)+a(9) (a(5)-6 a(6)+21 a(7)-56 a(8)+126 a(9)-252)+210=0 \]

OpenStudy (anonymous):

oh u solved for a0,a1,a2,... so nice thank u @eliassaab

OpenStudy (anonymous):

I got\[ \{\{a(0)\to 1,a(1)\to 5,a(2)\to 15,a(3)\to 30,a(4)\to 45,a(5)\to 51,a(6)\to 45,a(7)\to 30,a(8)\to 15,a(9)\to 5\}\} \] Which gives \[ P(x)= \]

OpenStudy (anonymous):

\[ P(x)= x^{10}+5 x^9+15 x^8+30 x^7+45 x^6+\\51 x^5+45 x^4+30 x^3+15 x^2+5 x+1=\left(x^2+x+1\right)^5 \]

OpenStudy (anonymous):

special thanks to MATHEMATICA and u so my solution is right

OpenStudy (anonymous):

I took a(10) =1

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

It seems the only polynomials of 2p degress that satisfy that P(x^2)=P(x)P(x-1) is of the from \[ \left ( 1+x+x^2 \right)^p \]

OpenStudy (anonymous):

The above statement is always true.